# The Very High Energy Afterglow of Structured Jets: GW 170817 and Prospects for Future Detections

### v1

@@ 5 Detectability of post-merger afterglows at VHE and prospects for the CTA
Another direct consequence of the impact of $n_{\mathrm{ext}}$ on SSC emission is that short GRB afterglows detected at VHE are more likely to originate from mergers occurring in a denser environment. Such VHE detections could therefore be unique probes to high-density media at the location of BNS mergers, which are sought clues for short merger time binaries. Fast mergers are indeed expected to occur close to their formation sites, where the external medium density is expected to be higher (see e.g. the discussion in Duque et al., 2020). Population synthesis studies have shown that the yet poorly-constrained phase of common envelope during binary stellar evolution can lead to the production of BNS with very short initial separations that merge rapidly, as discussed by e.g. Dominik et al. (2012). Favorable kick intensities and orientations during the supernovae could also tend to shrink the initial orbital separation and increase the orbital eccentricity for part of the BNS population (Kalogera, 1996). Observationally, the projected offset distribution of BNS mergers in their host galaxy seems to feature a tail below 1 half-light radius for $\sim 20\%$ of systems (Fong & Berger, 2013; Berger, 2014). These observations suffer from several observational biases like afterglow detection, host galaxy association and projection effects, probably leading to sample incompleteness, but give evidence for a population of fast mergers. If BNS mergers are the dominant site for r-process nucleosynthesis (see e.g. Lattimer & Schramm, 1974; Goriely et al., 2011; Just et al., 2015; Côté et al., 2017), another more indirect evidence is provided by the observed large scatter of [Eu/Fe] abundances in extremely metal-poor stars (see the compilation by Suda et al., 2008), which requires fast mergers allowing the production of r-process elements when the environment is still extremely metal-poor (see e.g. Vangioni et al., 2016; Côté et al., 2019; Dvorkin et al., 2021). It has already been suggested by Duque et al. (2020) that multi-messenger observations of BNS including the detection of the afterglow will be biased in favor of high-density environment, thus allowing to probe this potential population of fast mergers. We show here that future VHE afterglow detections either following the GW detection of a BNS merger, or the detection of a short GRB, can also directly probe this population.
@@ 6 Conclusion
We developed a detailed model of the afterglow of a laterally-structured jet from radio to VHE. Emission is produced in the shocked external medium behind the forward external shock and the dynamics does not include the late lateral expansion of the ejecta. Such effects, as well as the contribution of the reverse shock, will be implemented in the future. The main challenge addressed in this paper is the self-consistent calculation of the synchrotron and SSC emission while accounting for the two different IC regimes: Thomson and Klein-Nishina (KN). We based our approach on the method proposed by Nakar et al. (2009), extended to include additional effects: the maximum electron Lorentz factor $\gamma_{\mathrm{max}}$ is also computed self-consistently, leading to a realistic estimate of the high-energy cutoff of the synchrotron component, and the attenuation at high-energy due to pair production is also included. The values of $\gamma_{\mathrm{c}}$ and $Y(\gamma_{\mathrm{c}})$, as well as the spectral regimes are determined numerically. Synchrotron self-absorption, that is relevant at low radio frequencies, has not been included yet. The implementation of the model is computationally-efficient, allowing for afterglow fitting.
@@
We used this model to fit the multi-wavelength afterglow of GW 170817, including radio to X-ray observations up to 400 days and five early-time upper limits. We inferred the model parameters using three different assumptions for the emission: pure synchrotron, synchrotron and SSC assuming that all scatterings occur in Thomson regime, and self-consistent synchrotron and SSC calculation taking into account the KN regime. We obtain excellent fits in the three cases, with similar constraints on the parameters for the pure synchrotron case and the full calculation with SSC in KN regimes, whereas the case where all scatterings are assumed to be in the Thomson regime deviates significantly. The SSC emission in the afterglow of GW 170817 is indeed weak due to a strong KN attenuation, which confirms the need to include the self-consistent calculation of the SSC emission in the KN regime in afterglow models. The predicted VHE flux at the peak has a large dispersion but remains about two orders of magnitude below the upper limit obtained around the peak by H.E.S.S. This shows that the VHE afterglow of GW 170817 was undetectable by current instruments, but that future detections of similar events would break some degeneracies among the model parameters and allow better constraints on the physics of the relativistic ejecta.
@@
We then studied how such post-merger VHE afterglows may become detectable in the future, in the context of the improved sensitivity of the CTA. Beyond the evident effect of a less off-axis viewing angle, we show that the efficiency of the SSC emission becomes larger in external media with higher densities than GW 170817. For instance, a similar ultra-relativistic jet in a medium of density $\sim 1\,\mathrm{cm^{-3}}$ and seen slightly less off-axis than GW 170817 may become detectable by the CTA at 100-400 Mpc.
@@
This bias of VHE afterglow detections in favor of high-density environments offers a new probe of a possible population of fast mergers. With a short merger time, a merger should indeed occur in more central, and denser, regions of its host galaxy. Thus, future detections of such VHE afterglows would not only advance our physical understanding of relativistic jets produced after a BNS merger, but also constrain the physics of the evolution of massive binary systems.
Acknowledgements
@@
Clément Pellouin acknowledges funding support from the Initiative Physique des Infinis (IPI), a research training program of the Idex SUPER at Sorbonne University. The authors acknowledge the Centre National d’Études Spatiales (CNES) for financial support in this research project. This work made use of the Infinity computing cluster at IAP. The authors thank K. Mooley for providing GW 170817 afterglow observation data at github.com/kmooley/GW170817/. This work made use of Python 3.8 and of the Python packages arviz (Kumar et al., 2019), astropy (Astropy Collaboration et al., 2022), corner (Foreman-Mackey, 2016), emcee (Foreman-Mackey et al., 2019, 2013), matplotlib (Hunter, 2007), multiprocessing, numpy (Harris et al., 2020), pandas (McKinney, 2010), scipy (Virtanen et al., 2020).
@@ Appendix A Maximum electron Lorentz factor and synchrotron burnoff limit
The maximum Lorentz factor $\gamma_{\mathrm{max}}$ of accelerated electrons is defined in Section 2.3 and is reached when the acceleration timescale becomes longer than the cooling timescale, i.e.
@@
where the dynamical timescale $t_{\mathrm{dyn}}$ and the acceleration timescale $t_{\mathrm{acc}}$ are defined by Equations (20) and (28). From Equations (4243), the radiative timescale $t_{\mathrm{rad}}$ is given by
@@
In the present version of the model, we assume $\gamma_{\mathrm{max}}\gg\gamma_{\mathrm{m}}$ so that the calculation of $\gamma_{\mathrm{c}}$ and the identification of the radiative regime and associated normalized electron distribution $f(x)$ and Compton parameter $h(x)$ are independent of $\gamma_{\mathrm{max}}$ (see § 2.4.6). Therefore the maximum Lorentz factor is determined once this radiative regime is identified. We follow an iterative procedure where the segments of $h(x)$ are explored successively starting from the last segment $x>\widehat{x}_{1}$ (highest Lorentz factors):
@@
    We check if there is a solution of Equation (103) in this segment, with the usual approximation for $1+Y$, i.e.
@@
    where $Y(\gamma_{\mathrm{max}})$ is computed from Equation (79), which leads to
@@
    • If the next segment is still at least partially in fast cooling regime (i.e. the upper bound is above $\gamma_{\mathrm{c}}$), we start a new iteration at step 1. using this new segment;
    • otherwise we stop the iterative procedure and keep for $\gamma_{\mathrm{max}}$ the slow cooling solution $\gamma_{\mathrm{max}}=1/\widetilde{K}$.
In most cases, the maximum electron Lorentz factor is large enough to be in fast cooling regime ($\gamma_{\mathrm{max}}>\gamma_{\mathrm{c}}$ so that $t_{\mathrm{rad}}(\gamma_{\mathrm{max}})<t_{\mathrm{dyn}}$) with a negligible IC cooling ($Y(\gamma_{\mathrm{max}})<1$). In this case, the maximum Lorentz factor $\gamma_{\mathrm{max}}=\sqrt{{\gamma_{\rm c}^{\rm syn}}/{\widetilde{K}}}$ leads to the usual synchrotron burnoff limit (see e.g. Piran & Nakar 2010) for the high-frequency cutoff of the synchrotron spectrum in the comoving frame $h\nu_{\mathrm{max}}=h\nu_{\mathrm{syn}}(\gamma_{\mathrm{max}})=K_{2}B{\gamma_{\rm c}^{\rm syn}}/{\widetilde{K}}$:
@@
For the best-fit parameters of the afterglow of GW 170817 presented in Section 4, the Lorentz factor of the core jet at $t_{\mathrm{peak}}$ is $\Gamma\simeq 10$. The limit is thus expected at $\sim 1\leavevmode\nobreak\ \mathrm{GeV}$ in the observer frame, as seen in Fig. 9.
@@ Appendix B Radiative regimes and corresponding electron distributions
Figure 15: Definition of all possible cases for the radiative regime and the distribution of electrons in the plane $\gamma_{0}$ vs $\gamma_{\mathrm{self}}$. The Lorentz factors $\gamma_{\mathrm{m}}$ and $\gamma_{\mathrm{c}}$ are indicated by red and green solid lines either in the fast cooling regime with $\gamma_{\mathrm{m}}>\gamma_{\mathrm{c}}$ (left) or in the slow cooling regime with $\gamma_{\mathrm{m}}<\gamma_{\mathrm{c}}$ (right). The black solid line in the left panel corresponds to $\gamma_{0}=\gamma_{\mathrm{self}}$. The two Lorentz factors $\widehat{\gamma}_{\mathrm{m}}$ (red) and $\widehat{\gamma}_{\mathrm{c}}$ (green) are plotted in dashed lines. In some regions, additional Lorentz factors relevant to identify the radiative regime are also plotted: $\widetilde{\gamma}_{\mathrm{m}}$ (red dotted line) and $\widetilde{\gamma}_{\mathrm{c}}$ (green dotted line). The regions where there is no impact of the SSC cooling on the distribution of electrons are shaded in orange ($\widehat{\gamma}_{\mathrm{m,c}}<\gamma_{0}<\gamma_{\mathrm{c}}$) and blue ($\gamma_{0}<\widehat{\gamma}_{\mathrm{m,c}}$). This diagram allows to identify all possible orderings of $\gamma_{0}$, $\gamma_{\mathrm{m}}$, $\gamma_{\mathrm{c}}$, $\widehat{\gamma}_{\mathrm{m}}$, $\widehat{\gamma}_{\mathrm{m}}$ defining all possible radiative regimes. The corresponding limits for $\gamma_{\mathrm{self}}$ are indicated by vertical black dashed lines. All the possible cases are listed in Table 10 (fast cooling) and 11 (slow cooling). In practice, several cases have the same breaks and slopes for $f(x)$ and $h(x)$ and can therefore be treated simultaneously in the numerical implementation of the model. This allows to group $\left(F1,F2,F3,F4,F5\right)$, $\left(F8,F13,F14\right)$, $\left(F9,F15,F16\right)$, $\left(F10,F17\right)$, $\left(F20,F21\right)$, $\left(F22,F23\right)$, $\left(S1,S2,S3,S4,S5,S6,S7\right)$, $\left(S8,S9\right)$, $\left(S10,S12,S13\right)$ and $\left(S11,S14\right)$.
The IC cooling impacts the distribution of electrons if $\max{\left(\gamma_{\mathrm{c}};\tilde{\gamma}_{\mathrm{p}}\right)}<\gamma<\gamma_{0}$ (Equation (82)). As
@@
all possibles cases regarding the radiative regime and the distribution of electrons can be identified in the $\left(\gamma_{\mathrm{self}};\gamma_{0}\right)$ plane shown in Fig. 15. As an electron of Lorentz factor $\gamma$ can upscatter its own synchrotron photon only if $\gamma\leq\widehat{\gamma}$, i.e. $\gamma\leq\gamma_{\mathrm{self}}$, the x-axis quantifies the importance of the KN suppression: the lower $\gamma_{\mathrm{self}}$ (left side of the diagram), the more electrons are affected by KN effects. The y-axis quantifies the importance of IC cooling: the IC power of all electrons with $\gamma>\gamma_{0}$ is negligible compared to the synchrotron power. Therefore, the lower $\gamma_{0}$ (bottom of the diagram), the more electrons have a negligible IC cooling. We list in Tables 10 (fast cooling) and 11 (slow cooling) all these cases and provide the corresponding self-consistent solution (breaks and slopes) for the normalized distribution of electrons $f(x)$ (Equation (50)) and the normalized Compton parameter $h(x)$ (Equation (79)), as obtained following the detailed method described in Nakar et al. (2009). The break corresponding to the peak of the synchrotron spectrum is highlighted in red. We also provide in each case the expression of $\gamma_{0}$ obtained from the relation $Y(\gamma_{\mathrm{c}})=h(x_{\mathrm{c}})/h(x_{0})$ (Equation (84)), as needed in the iterative procedure to determine $Y(\gamma_{\mathrm{c}})$: see § 2.4.6.
The distribution of electrons is unaffected by the SSC cooling when $\gamma_{0}<\gamma_{\mathrm{c}}$ or $\gamma_{\mathrm{c}}<\gamma_{0}<\widehat{\gamma}_{\mathrm{m,c}}$. These two regions are shaded in orange and blue in Fig. 15. When $\gamma_{0}<\widehat{\gamma}_{\mathrm{m,c}}$ (cases F1 and S1), the IC cooling is negligible for all electrons and $\gamma_{0}$ is undefined. The solution is provided by the pure synchrotron case (§ 2.4.4). When $\widehat{\gamma}_{\mathrm{m,c}}<\gamma_{0}<\gamma_{\mathrm{c}}$ (cases F2 to F5 and S2 to S6), the normalized electron distribution $f(x)$ is also the same as in the pure synchrotron case but $\gamma_{0}$ is defined and its expression depends on the considered regime.
@@
Cases F24 and F25 are indeed more complex and require a specific treatment (see Section 3.2 Nakar et al. 2009), as a large number of breaks can appear in the distribution, that must be found iteratively. These two cases correspond to $\gamma_{\mathrm{c}}<\widehat{\gamma}_{\mathrm{m}}<\gamma_{\mathrm{self}}<\gamma_{\mathrm{m}}<\widehat{\gamma}_{\mathrm{c}}$. New breaks will appear in the spectral shape $f(x)$ every time that ${}^{(2n)}\widehat{\gamma}_{\mathrm{m}}=\widehat{\gamma}_{\mathrm{c}}$, where using Equation (107), we define the successive ${}^{(n)}\widehat{\gamma}$ as
@@
with $n\geq 1$. For $n=1$, the first value $\gamma_{\mathrm{self,cr,2}}=\left(\gamma_{\mathrm{m}}^{2}\gamma_{\mathrm{c}}\right)^{\frac{1}{3}}$ is plotted in Fig. 15 and is the frontier at low $\gamma_{\mathrm{self}}$ of the cases F24 and F25. On the other hand $\gamma_{\mathrm{self,cr,2n}}$ tends to $\gamma_{\mathrm{m}}$ when $n$ tends to infinity, i.e. the frontier at high $\gamma_{\mathrm{self}}$. Then, the values of $\gamma_{\mathrm{self,cr,2n}}$ for $n>1$ divide F24 and F25 in many sub-regions. We define the column $n$ by the condition $\gamma_{\mathrm{self,cr,2n}}<\gamma_{\mathrm{self}}<\gamma_{\mathrm{self,cr,2n+2}}$. In practice, $\gamma_{\mathrm{self,cr,2n}}$ rapidly converges, and the spectral shapes in these narrower columns rapidly become closer and closer to each other. This allows numerically to not distinguish columns above a certain $n_{\mathrm{max}}$. We use $n_{\mathrm{max}}=7$.
@@
101010 For each case, the first row of the table gives the list of breaks appearing either in the normalized electron distribution $f(x)$, i.e. the list of $x_{i}$, or in the normalized Compton parameter $h(x)$, i.e. list of relevant $\widehat{x}_{i}$, as well as all other necessary characteristic Lorentz factors to fully define the case in agreement with Fig. 15 (blue). The break corresponding to the peak of the synchrotron spectrum is highlighted in red. The second and third rows give the slopes $-\mathrm{d}\ln{f}/d\ln{x}$ and $-\mathrm{d}\ln{g}/d\ln{x}$ for each branch. Finally, the last column gives the expression of $\gamma_{0}$. Cases F24 and F25 require a specific treatment: see text in appendix B.
@@
Figure 16: Posterior joint and marginalized distributions of the model’s free parameters, $E_{\mathrm{0,iso}}^{\mathrm{c}}$, $\theta_{\mathrm{c}}$, $\theta_{\mathrm{v}}$, $n_{\mathrm{ext}}$, $\epsilon_{\mathrm{B}}$, $\epsilon_{\mathrm{e}}$, $\zeta$, $p$ $a$, $\Gamma_{0}^{\mathrm{c}}$, $b$ for the three fits of the afterglow of GW 170817 presented in section 4, which differ only by the treatment of the radiative processes: in blue, no SSC diffusions are taken into account (”no SSC”); in green, SSC diffusions are assumed to be only in the Thomson regime (”SSC (Thomson)”); in orange, the SSC diffusion is depleted at high energy in the Klein-Nishina regime (”SSC (with KN)”). The colored contours correspond to the $3\sigma$ confidence intervals for each model and each parameter.

### v2

@@ 5 Detectability of post-merger afterglows at very high energy, and prospects for the CTA
Another direct consequence of the impact of $n_{\mathrm{ext}}$ on SSC emission is that short GRB afterglows detected at VHE are more likely to originate from mergers that occur in a denser environment. These VHE detections could therefore be unique probes of high-density media at the location of BNS mergers, which are sought clues for binaries with short merger times. Fast mergers are indeed expected to occur close to their formation sites, where the external medium density is expected to be higher (see e.g. the discussion in Duque et al. 2020). Population synthesis studies have shown that the as yet poorly constrained common-envelope phase during the binary star evolution can lead to the production of BNS with very short initial separations that merge rapidly, as discussed for instance by Dominik et al. (2012). Favourable kick intensities and orientations during the supernovae could also tend to shrink the initial orbital separation and increase the orbital eccentricity for part of the BNS population (Kalogera 1996). Observationally, the projected offset distribution of BNS mergers in their host galaxy seems to feature a tail below one half-light radius for $\sim 20\%$ of the systems (Fong & Berger 2013; Berger 2014). These observations have several observational biases, such as afterglow detection, host galaxy association, and projection effects, which probably leads to sample incompleteness, but they give evidence for a population of fast mergers. If BNS mergers are the dominant site for r-process nucleosynthesis (see e.g. Lattimer & Schramm 1974; Goriely et al. 2011; Just et al. 2015; Côté et al. 2017), another more indirect evidence is provided by the observed large scatter of [Eu/Fe] abundances in extremely metal-poor stars (see the compilation by Suda et al. 2008), which requires fast mergers that allow the production of r-process elements when the environment is still extremely metal poor (see e.g. Vangioni et al. 2016; Côté et al. 2019; Dvorkin et al. 2021). It has already been suggested by Duque et al. (2020) that multi-messenger observations of BNS, including the detection of the afterglow, will be biased in favour of a high-density environment, thus allowing us to probe this potential population of fast mergers. We showed here that future VHE afterglow detections either following the GW detection of a BNS merger or the detection of a short GRB can also probe this population directly.
@@ 6 Conclusion
We developed a detailed model of the afterglow of a laterally structured jet from radio to VHE. Emission is produced in the shocked external medium behind the forward external shock, and the dynamics does not include the late lateral expansion of the ejecta. These effects, as well as the contribution of the reverse shock, will be implemented in the future. The main challenge addressed in this paper is the self-consistent calculation of the synchrotron and SSC emission while accounting for the two different IC regimes, the Thomson and KN regime. We based our approach on the method proposed by Nakar et al. (2009), which we extended to include additional effects: The maximum electron Lorentz factor $\gamma_{\mathrm{max}}$ was also computed self-consistently, leading to a realistic estimate of the high-energy cutoff of the synchrotron component, and the attenuation at high-energy due to pair production was included as well. The values of $\gamma_{\mathrm{c}}$ and $Y(\gamma_{\mathrm{c}})$ and of the spectral regimes were determined numerically. Synchrotron self-absorption, which is relevant at low radio frequencies, is not yet included. The implementation of the model is computationally efficient and allows a fitting of afterglow data.
@@
We used this model to fit the multi-wavelength afterglow of GW 170817, including radio to X-ray observations up to 400 days and five early-time upper limits. We inferred the model parameters using three different assumptions for the emission: (i) Purely synchrotron, (ii) synchrotron and SSC, assuming that scatterings occur in Thomson regime, and (iii) self-consistent synchrotron and SSC calculation taking the KN regime into account. We obtained excellent fits in the three cases, with similar constraints on the parameters for the purely synchrotron case and the full calculation with SSC in KN regimes, whereas the case in which scattering was assumed to be in the Thomson regime deviated significantly. The SSC emission in the afterglow of GW 170817 is indeed weak due to the strong KN attenuation, which confirms the need to include the self-consistent calculation of the SSC emission in the KN regime in afterglow models. The predicted VHE flux at the peak has a large dispersion, but remains about two orders of magnitude below the upper limit obtained around the peak by H.E.S.S. This shows that the VHE afterglow of GW 170817 was undetectable by current instruments, but that future detections of similar events would break some degeneracies among the model parameters and would provide better constraints on the physics of the relativistic ejecta.
@@
We then studied how these post-merger VHE afterglows may become detectable in the future in the context of the improved sensitivity of the CTA. Beyond the evident effect of a less off-axis viewing angle, we showed that the efficiency of the SSC emission becomes higher in external media with higher densities than GW 170817. For instance, a similar ultra-relativistic jet in a medium with a density of $\sim 1\,\mathrm{cm^{-3}}$ and seen slightly less off-axis than GW 170817 may become detectable by the CTA at 100-400 Mpc.
@@
This bias of VHE afterglow detections in favour of high-density environments offers a new probe of a possible population of fast mergers. With a short merger time, a merger should indeed occur in more central, and denser, regions of its host galaxy. Thus, future detections of these VHE afterglows would not only advance our physical understanding of relativistic jets that are produced after a BNS merger, but would also constrain the physics of the evolution of massive binary systems.
Acknowledgements.
@@
Clément Pellouin acknowledges funding support from the Initiative Physique des Infinis (IPI), a research training program of the Idex SUPER at Sorbonne University. The authors acknowledge the Centre National d’Études Spatiales (CNES) for financial support in this research project. This work made use of the Infinity computing cluster at IAP. The authors thank K. Mooley for providing GW 170817 afterglow observation data at github.com/kmooley/GW170817/. This work made use of Python 3.8 and of the Python packages arviz (Kumar et al. 2019), astropy (Astropy Collaboration et al. 2022), corner (Foreman-Mackey 2016), emcee (Foreman-Mackey et al. 2019, 2013), matplotlib (Hunter 2007), multiprocessing, numpy (Harris et al. 2020), pandas (McKinney 2010), scipy (Virtanen et al. 2020).
@@ Appendix A Maximum electron Lorentz factor and synchrotron burnoff limit
The maximum Lorentz factor $\gamma_{\mathrm{max}}$ of accelerated electrons is defined in Sect. 2.3 and is reached when the acceleration timescale becomes longer than the cooling timescale,
@@
where the dynamical timescale $t_{\mathrm{dyn}}$ and the acceleration timescale $t_{\mathrm{acc}}$ are defined by Eqs. (20) and (28). From Eqs. (42) and (43), the radiative timescale $t_{\mathrm{rad}}$ is given by
@@
In the current version of the model, we assume $\gamma_{\mathrm{max}}\gg\gamma_{\mathrm{m}}$ so that the calculation of $\gamma_{\mathrm{c}}$ and the identification of the radiative regime and associated normalised electron distribution $f(x)$ and Compton parameter $h(x)$ are independent of $\gamma_{\mathrm{max}}$ (see Sect. 2.4.6). Therefore the maximum Lorentz factor is determined once this radiative regime is identified. We follow an iterative procedure where the segments of $h(x)$ are explored successively starting from the last segment $x>\widehat{x}_{1}$ (highest Lorentz factors):
@@
    We check if there is a solution of Eq. (103) in this segment, with the usual approximation for $1+Y$,
@@
    where $Y(\gamma_{\mathrm{max}})$ is computed from Eq. (79), which leads to
@@
    • If the next segment is still at least partially in fast-cooling regime (i.e. the upper bound is above $\gamma_{\mathrm{c}}$), we start a new iteration at step 1. using this new segment;
    • otherwise we stop the iterative procedure and keep for $\gamma_{\mathrm{max}}$ the slow-cooling solution $\gamma_{\mathrm{max}}=1/\widetilde{K}$.
In most cases, the maximum electron Lorentz factor is large enough to be in fast-cooling regime ($\gamma_{\mathrm{max}}>\gamma_{\mathrm{c}}$ so that $t_{\mathrm{rad}}(\gamma_{\mathrm{max}})<t_{\mathrm{dyn}}$) with a negligible IC cooling ($Y(\gamma_{\mathrm{max}})<1$). In this case, the maximum Lorentz factor $\gamma_{\mathrm{max}}=\sqrt{{\gamma_{\rm c}^{\rm syn}}/{\widetilde{K}}}$ leads to the usual synchrotron burnoff limit (see e.g. Piran & Nakar 2010) for the high-frequency cutoff of the synchrotron spectrum in the comoving frame $h\nu_{\mathrm{max}}=h\nu_{\mathrm{syn}}(\gamma_{\mathrm{max}})=K_{2}B{\gamma_{\rm c}^{\rm syn}}/{\widetilde{K}}$:
@@
For the best-fit parameters of the afterglow of GW 170817 presented in Sect. 4, the Lorentz factor of the core jet at $t_{\mathrm{peak}}$ is $\Gamma\simeq 10$. The limit is thus expected at $\sim 1\leavevmode\nobreak\ \mathrm{GeV}$ in the observer frame, as seen in Fig. 9.
@@ Appendix B Radiative regimes and corresponding electron distributions
Figure 15: Definition of all possible cases for the radiative regime and the distribution of electrons in the plane $\gamma_{0}$ vs $\gamma_{\mathrm{self}}$. The Lorentz factors $\gamma_{\mathrm{m}}$ and $\gamma_{\mathrm{c}}$ are indicated by red and green solid lines either in the fast-cooling regime with $\gamma_{\mathrm{m}}>\gamma_{\mathrm{c}}$ (left) or in the slow-cooling regime with $\gamma_{\mathrm{m}}<\gamma_{\mathrm{c}}$ (right). The black solid line in the left panel corresponds to $\gamma_{0}=\gamma_{\mathrm{self}}$. The two Lorentz factors $\widehat{\gamma}_{\mathrm{m}}$ (red) and $\widehat{\gamma}_{\mathrm{c}}$ (green) are plotted in dashed lines. In some regions, additional Lorentz factors relevant to identify the radiative regime are also plotted: $\widetilde{\gamma}_{\mathrm{m}}$ (red dotted line) and $\widetilde{\gamma}_{\mathrm{c}}$ (green dotted line). The regions where there is no impact of the SSC cooling on the distribution of electrons are shaded in orange ($\widehat{\gamma}_{\mathrm{m,c}}<\gamma_{0}<\gamma_{\mathrm{c}}$) and blue ($\gamma_{0}<\widehat{\gamma}_{\mathrm{m,c}}$). This diagram allows to identify all possible orderings of $\gamma_{0}$, $\gamma_{\mathrm{m}}$, $\gamma_{\mathrm{c}}$, $\widehat{\gamma}_{\mathrm{m}}$, $\widehat{\gamma}_{\mathrm{m}}$ defining all possible radiative regimes. The corresponding limits for $\gamma_{\mathrm{self}}$ are indicated by vertical black dashed lines. All the possible cases are listed in Table 10 (fast cooling) and 11 (slow cooling). In practice, several cases have the same breaks and slopes for $f(x)$ and $h(x)$ and can therefore be treated simultaneously in the numerical implementation of the model. This allows to group $\left(F1,F2,F3,F4,F5\right)$, $\left(F8,F13,F14\right)$, $\left(F9,F15,F16\right)$, $\left(F10,F17\right)$, $\left(F20,F21\right)$, $\left(F22,F23\right)$, $\left(S1,S2,S3,S4,S5,S6,S7\right)$, $\left(S8,S9\right)$, $\left(S10,S12,S13\right)$ and $\left(S11,S14\right)$.
The IC cooling impacts the distribution of electrons if $\max{\left(\gamma_{\mathrm{c}};\tilde{\gamma}_{\mathrm{p}}\right)}<\gamma<\gamma_{0}$ (Eq. (82)). As
@@
all possibles cases regarding the radiative regime and the distribution of electrons can be identified in the $\left(\gamma_{\mathrm{self}};\gamma_{0}\right)$ plane shown in Fig. 15. As an electron of Lorentz factor $\gamma$ can upscatter its own synchrotron photon only if $\gamma\leq\widehat{\gamma}$, that is, if $\gamma\leq\gamma_{\mathrm{self}}$, the x-axis quantifies the importance of the KN suppression: the lower $\gamma_{\mathrm{self}}$ (left side of the diagram), the more electrons are affected by KN effects. The y-axis quantifies the importance of IC cooling: the IC power of all electrons with $\gamma>\gamma_{0}$ is negligible compared to the synchrotron power. Therefore, the lower $\gamma_{0}$ (bottom of the diagram), the more electrons have a negligible IC cooling. We list in Tables 10 (fast cooling) and 11 (slow cooling) all these cases and provide the corresponding self-consistent solution (breaks and slopes) for the normalised distribution of electrons $f(x)$ (Eq. (50)) and the normalised Compton parameter $h(x)$ (Eq. (79)), as obtained following the detailed method described in Nakar et al. (2009). The break corresponding to the peak of the synchrotron spectrum is highlighted in red. We also provide in each case the expression of $\gamma_{0}$ obtained from the relation $Y(\gamma_{\mathrm{c}})=h(x_{\mathrm{c}})/h(x_{0})$ (Eq. (84)), as needed in the iterative procedure to determine $Y(\gamma_{\mathrm{c}})$: see Sect. 2.4.6.
The distribution of electrons is unaffected by the SSC cooling when $\gamma_{0}<\gamma_{\mathrm{c}}$ or $\gamma_{\mathrm{c}}<\gamma_{0}<\widehat{\gamma}_{\mathrm{m,c}}$. These two regions are shaded in orange and blue in Fig. 15. When $\gamma_{0}<\widehat{\gamma}_{\mathrm{m,c}}$ (cases F1 and S1), the IC cooling is negligible for all electrons and $\gamma_{0}$ is undefined. The solution is provided by the purely synchrotron case (Sect. 2.4.4). When $\widehat{\gamma}_{\mathrm{m,c}}<\gamma_{0}<\gamma_{\mathrm{c}}$ (cases F2 to F5 and S2 to S6), the normalised electron distribution $f(x)$ is also the same as in the purely synchrotron case but $\gamma_{0}$ is defined and its expression depends on the considered regime.
@@
Cases F24 and F25 are indeed more complex and require a specific treatment (see Sect. 3.2 Nakar et al. 2009), as a large number of breaks can appear in the distribution, that must be found iteratively. These two cases correspond to $\gamma_{\mathrm{c}}<\widehat{\gamma}_{\mathrm{m}}<\gamma_{\mathrm{self}}<\gamma_{\mathrm{m}}<\widehat{\gamma}_{\mathrm{c}}$. New breaks will appear in the spectral shape $f(x)$ every time that ${}^{(2n)}\widehat{\gamma}_{\mathrm{m}}=\widehat{\gamma}_{\mathrm{c}}$, where using Eq. (107), we define the successive ${}^{(n)}\widehat{\gamma}$ as
@@
with $n\geq 1$. For $n=1$, the first value $\gamma_{\mathrm{self,cr,2}}=\left(\gamma_{\mathrm{m}}^{2}\gamma_{\mathrm{c}}\right)^{\frac{1}{3}}$ is plotted in Fig. 15 and is the frontier at low $\gamma_{\mathrm{self}}$ of the cases F24 and F25. On the other hand $\gamma_{\mathrm{self,cr,2n}}$ tends to $\gamma_{\mathrm{m}}$ when $n$ tends to infinity, that is, the frontier at high $\gamma_{\mathrm{self}}$. Then, the values of $\gamma_{\mathrm{self,cr,2n}}$ for $n>1$ divide F24 and F25 in many sub-regions. We define the column $n$ by the condition $\gamma_{\mathrm{self,cr,2n}}<\gamma_{\mathrm{self}}<\gamma_{\mathrm{self,cr,2n+2}}$. In practice, $\gamma_{\mathrm{self,cr,2n}}$ rapidly converges, and the spectral shapes in these narrower columns rapidly become closer and closer to each other. This allows numerically to not distinguish columns above a certain $n_{\mathrm{max}}$. We use $n_{\mathrm{max}}=7$.
@@
101010 For each case, the first row of the table gives the list of breaks appearing either in the normalised electron distribution $f(x)$ (i.e. the list of $x_{i}$), or in the normalised Compton parameter $h(x)$ (i.e. the list of relevant $\widehat{x}_{i}$), as well as all other necessary characteristic Lorentz factors to fully define the case in agreement with Fig. 15 (blue). The break corresponding to the peak of the synchrotron spectrum is highlighted in red. The second and third rows give the slopes $-\mathrm{d}\ln{f}/d\ln{x}$ and $-\mathrm{d}\ln{g}/d\ln{x}$ for each branch. Finally, the last column gives the expression of $\gamma_{0}$. Cases F24 and F25 require a specific treatment: see text in appendix B.